# The foot, the fan, and the cuprate phase diagram: Fermi-volume-changing quantum phase transitions

### v1

@@ 4 The ‘fan’: QPT without symmetry breaking
In principle, the symmetry breaking theory for the ‘foot’ in (7) can be extended to higher temperatures to also provide a theory for the ‘fan’, and this may well be the appropriate approach for many correlated electron compounds [80]. The theory of the ‘fan’ in Refs. [39, 47] using the 2dYSYK model in (12) applies to the extended bosons in Fig. 6.
@@
However, such an approach does not apply to the hole-doped cuprates in particular, because the spin correlation length in the pseudogap regime is too small to explain the large gap seen in photoemission in the anti-nodal region of the Brillouin zone [81]. Therefore, as we noted in Section 2, we model the pseudogap metal of the hole-doped cuprates as a fractionalized Fermi liquid (FL*), in which the electrons missing from the Fermi surface count are placed in a spin liquid with fractionalization [57, 58].
@@
Figure 7: RIXS spectrum of La2-xSrxCuO4 from Ref. [82]. Note how the sharp spin wave excitation at $x=0$ turns into broad continua at non-zero $x$. The latter have been argued [83] to be the spinons needed for the existence of the FL* state.
@@
A FL* phase is required to have a background spin liquid, and we can ask for experimental evidence for the spinons of such a spin liquid. We have argued recently [83] that such evidence is readily available in existing observations: spinons are the most natural interpretation of RIXS measurements of the higher energy spin fluctuation spectrum in the cuprates [84, 82]. As shown in Fig. 7, the doped cuprates have a broad continuum of high energy spin excitations, which become sharp spin waves in the zero doping limit. These have been labeled “intense paramagnons” [84], but this appears implausible because of the absence of a large Fermi surface in the underdoped regime. A spin wave interpretation is also not tenable in the absence of antiferromagnetic order. Instead the observations of Ref. [82] are similar to those in the triangular lattice antiferromagnet KYbSe2 [85]; the latter displays a continuum at high energies which has been modeled in a theory of spinons [86, 85], along with weak antiferromagnetic order at low temperatures.
@@
Evidence for a FL* phase at low doping has also emerged in quantum simulations using ultracold atoms [87, 88], which show a clear transformation from a ‘polaronic metal’ (which we identify as FL*) to a FL with increasing doping. Numerical computations on ancilla wavefunctions of FL* [89, 90] agree well with the observations.
@@
Finally, as we have shown recently [51] (see discussion below (11)), the transition from FL* to FL has a unique signature in having a singular enhancement of thermopower in a ‘skewed marginal Fermi liquid’ [91] in the presence of spatial disorder, and this is consistent with observations in the hole-doped cuprates [92, 93].
@@ 4.1 Theory of the FL* phase
Figure 8: Phases of a Kondo lattice model with $f$-electron spins coupled to a conduction band of $c$ electrons of density $p$. At large Kondo coupling, $J_{K}$, the condensation of a Higgs boson $\Phi$ leads to the conventional FL which obeys the Luttinger count. The FL* phase is obtained when $\Phi$ is uncondensed, and the $f$ electrons form a spin liquid whose gauge symmetry is preserved.
@@
Obtaining a FL* phase in a single band model is somewhat more subtle, and we illustrate the ancilla approach [94, 95] in Fig. 9.
@@
Photoemission spectra obtained by this approach have been successfully compared with experiments in Ref. [56] (one such result was shown in Fig. 4). For our purposes, it is sufficient to note that the FL* to FL transition is also driven by the condensation of a Higgs boson $\Phi$, but the transition is now ‘inverted’ with respect to the Kondo lattice model, as shown in Fig. 10.
@@ 4.2 FL-FL* QPT
We now address the nature of the FL-FL* QPT in Fig. 4, in the presence of spatial disorder. We will do this for the case of the Kondo lattice transition in Fig. 8 only, as the results for the single-band model of Fig. 9 are quite similar [52, 51, 94, 95, 96]. The theory is similar in structure to that for the symmetry breaking case in (7), with the Higgs field $\Phi$ replacing the SDW order parameter $\phi$: we have the imaginary time Lagrangian
@@
where we have omitted the emergent gauge field under which the Higgs field $\Phi$ and the spinons $f$ are charged [58]. It is remarkable that the critical properties of the QPT without symmetry breaking described by $\mathcal{L}_{2}$ are essentially the same as those of the QPT with symmetry-breaking described by $\mathcal{L}_{1}$ in (7): hence the claim of a universal theory of strange metals in Ref. [39]. There is, however, one important difference between $\mathcal{L}_{1}$ and $\mathcal{L}_{2}$. The field $\Phi$ is complex, and there is a pronounced particle-hole asymmetry in $\mathcal{L}_{2}$, which is absent for $\mathcal{L}_{1}$ [51]. Specifically, in the regime of extended boson eigenstates (see Fig. 6), the renormalized propagator for the boson has the form [35, 51]
$D({\bm{q}},i\Omega)\sim\frac{1}{Kq^{2}+c_{1}|\Omega|-ic_{2}\Omega+m^{2}(T)}$ (11)
@@
Figure 11: The top panels display the measurements of transport relaxation time, $\tau$, obtained from the optical conductivity in Ref. [19]. The bottom panels show computation of the same quantity in the 2dYSYK model in Ref. [47].
We are interested in applying $\mathcal{L}_{2}$ to the ‘fan’ at higher temperatures, and so we can focus on the case where all the fermionic and bosonic eigenmodes are extended. Here, it is important [39] to apply a mapping which transfers the ‘random mass’ disorder $\delta s({\bm{r}})$ to a random Yukawa coupling disorder $g^{\prime}({\bm{r}})$ by a space dependent rescaling of $\Phi$. This has the salutary effect of accounting for the leading effect of the random spatial structure of the bosonic eigenmodes, and of the Harris disorder, and placing them in a form amenable to methods used in the exact solution of the SYK model [39]. In this manner we obtain the 2dYSYK theory:
@@
The coupling $g^{\prime}({\bm{r}})$ obeys $\overline{g^{\prime}({\bm{r}})}=0$, $\overline{g^{\prime}({\bm{r}})g^{\prime}({\bm{r}^{\prime}})}=g^{\prime 2}\delta({\bm{r}}-{\bm{r}^{\prime}})$. The theory $\mathcal{L}_{3}$ has been studied perturbatively in $g$ and $\color[rgb]{.75,0,.25}g^{\prime}$ in Ref. [39], and in a fully self-consistent SYK-type solution for $g=0$ in Refs. [35, 47, 51]. The results display marginal Fermi liquid behavior in the fermion Green’s function, and linear-in-$T$ resistivity. Importantly, the Planckian behavior of the relaxation time of the SYK model in (1) is preserved (see Fig. 11); but the zero temperature entropy of the SYK model in (2) is not, and the 2dYSYK model displays a $\sim T\ln(1/T)$ entropy at the critical point, again consistent with observations.
@@ 5 Confinement crossover
We turn to the confinement crossover in Fig. 4, from the FL* state at higher temperature, to the symmetry-broken states without fractionalization at low temperatures. A framework for addressing this crossover was presented in Ref. [97]. The key ingredient is a SU(2) fundamental Higgs field $B$ (this is distinct from the Higgs field $\Phi$ in Section 4, which is assumed condensed in the underdoped region). The SU(2) gauge field arises from the underlying $\pi$-flux spin liquid [98] of the FL* state. The Higgs field $B$ also carries an electromagnetic charge $+e$. We assume that the pseudogap has $B$ uncondensed, and the low temperature confining phases appear via the condensation of $B$. A number of puzzles on the cuprate phase diagram can be addressed by this method:
@@
  • 5.
    The confinement crossover from FL* to the stripy SDW state at low temperatures remains to be understood [109]. It involves a delicate interplay between the charge order associated with the condensation of $B$, and the confinement of the fermionic spinons via the SU(2) gauge field [110]. Insights from numerical studies [111] should help understand this crossover.

### v5

@@ 4 The ‘fan’: QPT without symmetry breaking
In principle, the symmetry breaking theory for the ‘foot’ in (7) can be extended to higher temperatures to also provide a theory for the ‘fan’, and this may well be the appropriate approach for many correlated electron compounds [107]. The theory of the ‘fan’ in Refs. [43, 51] using the 2dYSYK model in (12) below also applies to the extended SDW bosons in Fig. 7.
@@
However, such an approach does not apply to the hole-doped cuprates in particular, because the spin correlation length in the pseudogap regime is too small to explain the large gap seen in photoemission in the anti-nodal region of the Brillouin zone [108, 73]. Therefore, as discussed in Section 2, we model the pseudogap metal of the hole-doped cuprates by a metallic state in which the electrons missing from the Fermi surface count are placed in a spin liquid with fractionalization [68, 69, 76, 77]. This leads to a FL-FL* QPT description of the higher temperature quantum critical fan, which we discuss in the present section. Further discussion of the nature of the pseudogap metal itself is deferred to Section 5.
@@
Figure 8: RIXS spectrum of La2-xSrxCuO4 from Robarts et al. [109]. Note how the sharp spin wave excitation at $x=0$ turns into broad continua at non-zero $x$. The latter have been argued by Bonetti et al. [110] to be the spinons needed for the theory of the pseudogap metal. See also the comparison with numerics in Fig. 9.
@@
We can ask for experimental evidence for the spinons of the spin liquid underlying the pseudogap metal. We have argued recently [110] that such evidence is readily available in existing observations: spinons are the most natural interpretation of RIXS measurements of the higher energy spin fluctuation spectrum in the cuprates [111, 109]. As shown in Fig. 8, the doped cuprates have a broad continuum of high energy spin excitations, which become sharp spin waves in the zero doping limit. These continuum excitations have been labeled “intense paramagnons” [111], but this appears implausible because of the absence of a large Fermi surface in the underdoped regime. A spin wave interpretation is also not tenable in the absence of antiferromagnetic order. We also note the recent NMR observations of a spin gap which have been associated with ‘short-range spin singlets’ [112].
Figure 9: Dynamic structure factor of the square lattice Heisenberg model with first ($J_{1}$) and second ($J_{2}$) neighbor exchange interactions from Ferrari and Becca [113], $J_{2}/J_{1}=0$ (upper left), 0.3 (upper right), 0.45 (lower left), and 0.55 (lower right). Note the evolution from sharp spin-wave modes at $J_{2}/J_{1}=0$ to spinon continua at larger $J_{2}/J_{1}$. Compare to the evolution of the cuprate spectrum in Fig. 8 with increasing doping between momentum $q=(0,0)$ and $q=(0,\pi)$ and between $q=(0,0)$ and $q=(\pi/2,\pi/2)$.
Instead the observations of Robarts et al. [109] are naturally interpreted as spinons. Note the evolution of their observations with increasing doping from a sharp spin-wave spectrum in Figs. 8a,b to a broad continuum in Figs. 8c-f. This matches the trends in the evolution of the dynamic spin structure factor of the insulating square lattice antiferromagnet with increasing second neighbor exchange, as computed by Ferrari and Becca [113] using fermionic spinons, and shown in Fig. 9. Indeed, the cuprate observations are the analog of those in the triangular lattice antiferromagnet KYbSe2 [61]. As discussed in Section 2 and Fig. 5, KYbSe2 displays a continuum at high energies which has been modeled in a theory of bosonic spinons [62, 61], along with weak antiferromagnetic order at low temperatures.
@@
We recall evidence via photoemission [9] and ADMR [114] for small pocket Fermi surfaces in the pseudogap regime of the hole-doped cuprates. Evidence for small Fermi surfaces at low doping has also emerged in quantum simulations using ultracold atoms [115, 116], which show a clear transformation from a ‘polaronic metal’ to a FL with increasing doping. Numerical computations on ancilla wavefunctions of FL* [117, 118] agree well with these observations.
@@
Finally, as we have shown recently [56] (see discussion below (11)), the transition from FL to FL* has a unique signature in having a singular enhancement of thermopower in a ‘skewed marginal Fermi liquid’ [119] in the presence of spatial disorder, and this is consistent with observations in the hole-doped cuprates [120, 121].
@@ 4.1 Theory of the FL* phase
Figure 10: Phases of a Kondo lattice model with $f$-electron spins coupled to a conduction band of $c$ electrons of density $p$. At large Kondo coupling, $J_{K}$, the condensation of a Higgs boson $\Phi$ leads to the conventional FL which obeys the Luttinger count. The FL* phase is obtained when $\Phi$ is uncondensed, and the $f$ electrons form a spin liquid whose gauge symmetry is preserved. Maksimovic et al. [122] present evidence for such a transition in CeCoIn5.
@@
Obtaining a FL* phase in a single band model is somewhat more subtle [123, 124, 125] (see Fig. 14), and we illustrate the ancilla approach [126, 127, 128] in Fig. 11.
@@
Photoemission spectra obtained by this approach have been successfully compared with experiments in Ref. [129]. For our purposes, it is sufficient to note that the FL to FL* transition is also driven by the condensation of a Higgs boson $\Phi$, but the transition is now ‘inverted’ with respect to the Kondo lattice model, as shown in Fig. 12.
@@ 4.2 FL-FL* QPT
We now address the nature of the FL-FL* QPT in Fig. 4, in the presence of spatial disorder. We will do this for the case of the Kondo lattice transition in Fig. 10 only, as the results for the single-band model of Fig. 11 differ only in their gauge structure [57, 56, 126, 127, 130], and this is unimportant at the level of the SYK-type equations in (13). The theory is similar in structure to that for the symmetry breaking case in (7), with the Higgs field $\Phi$ replacing the SDW order parameter $\phi$: we have the imaginary time Lagrangian
@@
where we have omitted the emergent gauge field under which the Higgs field $\Phi$ and the spinons $f$ are charged [69]. It is remarkable that, in the presence of spatial disorder, the critical properties of the QPT without symmetry breaking described by $\mathcal{L}_{2}$ are essentially the same as those of the QPT with symmetry-breaking described by $\mathcal{L}_{1}$ in (7): hence the claim of a universal theory of strange metals in Ref. [43]. There is, however, one important difference between $\mathcal{L}_{1}$ and $\mathcal{L}_{2}$. The Higgs field $\Phi$ is complex, and there is a pronounced particle-hole asymmetry in $\mathcal{L}_{2}$, which is absent for $\mathcal{L}_{1}$ [56]. Specifically, in the regime of extended boson eigenstates (see Fig. 7), the renormalized propagator for the Higgs boson has the form [39, 56]
$D(i\Omega,{\bm{q}})\sim\frac{1}{K{\bm{q}}^{2}+c_{1}|\Omega|-ic_{2}\Omega+m^{2}(T)}$ (11)
@@
Figure 13: The top panels display the measurements of transport relaxation time, $\tau$, obtained from the optical conductivity in Michon et al. [23]. The bottom panels show computation of the same quantity in the 2dYSYK model by Li et al. [51].
We are interested in applying $\mathcal{L}_{2}$ to the ‘fan’ at higher temperatures, and so we can focus on the case where all the fermionic and bosonic eigenmodes are extended. However, even in the extended boson regime, a perturbative averaging over the disorder $\delta s({\bm{r}})$ in (10) does not properly account for its strong effects, and is not the correct way to proceed. The key step [43] is to apply a mapping which transfers the ‘random mass’ disorder $\delta s({\bm{r}})$ to a random Yukawa coupling disorder $g^{\prime}({\bm{r}})$ by a space dependent rescaling of $\Phi$. This has the salutary effect of accounting exactly for the leading effect of the random spatial structure of the bosonic eigenmodes, and of the Harris disorder, and placing the randomness in a form amenable to methods used in the exact solution of the SYK model [43]. In this manner we obtain the 2dYSYK theory:
@@
The coupling $g^{\prime}({\bm{r}})$ obeys $\overline{g^{\prime}({\bm{r}})}=0$, $\overline{g^{\prime}({\bm{r}})g^{\prime}({\bm{r}^{\prime}})}=g^{\prime 2}\delta({\bm{r}}-{\bm{r}^{\prime}})$. The theory $\mathcal{L}_{3}$ has been studied perturbatively in $g$ and $\color[rgb]{.75,0,.25}g^{\prime}$ in Ref. [43]. After adding a large number of flavors and making the couplings random in flavor space, as in the SYK model, we can obtain the fully self-consistent solution SYK-type equations [43]:
$\Sigma(\tau,{\bm{r}})=g^{2}D(\tau,{\bm{r}})G(\tau,{\bm{r}})+{\color[rgb]{.75,0,.25}v^{2}\delta^{2}({\bm{r}})}G(\tau,{\bm{r}})+{\color[rgb]{.75,0,.25}{g^{\prime}}^{2}\delta^{2}({\bm{r}})}G(\tau,{\bm{r}})D(\tau,{\bm{r}}),$
$\Pi(\tau,{\bm{r}})=-g^{2}G(-\tau,-{\bm{r}})G(\tau,{\bm{r}})-{\color[rgb]{.75,0,.25}{g^{\prime}}^{2}G(-\tau,{\bm{r}})\delta^{2}({\bm{r}})}G(\tau,{\bm{r}}),$
$G(i\omega,{\bm{k}})=\frac{1}{i\omega-\varepsilon({\bm{k}})-\Sigma(i\omega,{\bm{k}})},$
$D(i\Omega,{\bm{q}})=\frac{1}{\Omega^{2}/c^{2}+K{\bm{q}}^{2}+s-\Pi(i\Omega,{\bm{q}})}\,,$ (13)
where $G$ is the electron Green’s function, $D$ is the boson Green’s function, and $\Sigma$ and $\Pi$ are the corresponding self energies. The equations have been written down for the simpler case where $\varepsilon({\bm{k}})=\varepsilon_{1}({\bm{k}})$, which applies to the SDW case in (7); in the presence of elastic scattering from $v({\bm{r}})$, the solutions are insensitive to the precise Fermi surfaces. The solution of (12) by (13) should be valid as long as the $\Phi$ eigenmodes remain extended. The equations (13) have been solved at $g=0$ in Refs. [39, 41, 43, 51, 56]. The solutions display marginal Fermi liquid behavior in the fermion Green’s function, and linear-in-$T$ resistivity. Importantly, the Planckian behavior of the relaxation time of the SYK model in (1) is preserved (see Fig. 13); but the zero temperature entropy of the SYK model in (2) is not. Instead, the 2dYSYK model solution via (13) displays a $\sim T\ln(1/T)$ entropy at the critical point, again consistent with observations.
A theory closely related to (12) has also been studied using non-pertubative quantum Monte Carlo at $g=0$ in Ref. [54]. Their results at low $T$ display a localized boson foot regime, similar to that of the random mass model (8) in Ref. [49]. This agreement is evidence for the validity of the mapping from (10) to (12).
@@ 5 Confinement crossover
Finally, we turn to the confinement crossover in Fig. 4, from the pseudogap metal state at higher temperature, to the symmetry-broken states without fractionalization at low temperatures.
We need to begin by discussing the nature of the pseudogap metal itself, and in particular the nature of the underlying spin liquid. An important subtlety is that the square lattice spin liquid has dual representations employing fermionic or bosonic spinons.
Christos et al. [131] have proposed the $\pi$-flux spin liquid with Dirac fermion spinons [132]. The fermionic spinons $f_{i\sigma}$ are obtained by transforming the electrons $c_{i\sigma}$ to a rotating reference frame in pseudospin space [133]
$\left(\begin{array}[]{cc}c_{i\uparrow}&-c_{i\downarrow}^{\dagger}\\ c_{i\downarrow}&c_{i\uparrow}^{\dagger}\end{array}\right)\sim\left(\begin{array}[]{cc}f_{i\uparrow}&-f_{i\downarrow}^{\dagger}\\ f_{i\downarrow}&f_{i\uparrow}^{\dagger}\end{array}\right)\left(\begin{array}[]{cc}B_{i1}^{\ast}&B_{i2}\\ -B_{i2}^{\ast}&B_{i1}\end{array}\right)\,,$ (20)
where $(B_{i1},B_{i2})$ are charge $e$ spinless bosons. This description has an emergent SU(2) gauge field, and $B$ is a SU(2) fundamental Higgs field [134]. This approach has been developed for a theory of the pseudogap as a FL* with a $\pi$-flux spin liquid, and its lower $T$ confinement transitions in Refs. [131, 135, 136].
Using fermion-boson duality, the $\pi$-flux spin liquid was argued by Wang et al. [137] to have dual description as the $\mathbb{CP}^{1}$ theory of bosonic spinons [138]. This bosonic spinon description is obtained by the complementary representation of fractionalization by a rotating reference frame in spin space [72, 73, 74, 75]
$\left(\begin{array}[]{cc}c_{i\uparrow}&-c_{i\downarrow}^{\dagger}\\ c_{i\downarrow}&c_{i\uparrow}^{\dagger}\end{array}\right)\sim\left(\begin{array}[]{cc}z_{i\uparrow}&-z_{i\downarrow}^{\ast}\\ z_{i\downarrow}&z_{i\uparrow}^{\ast}\end{array}\right)\left(\begin{array}[]{cc}\psi_{i+}&-\psi_{i-}^{\dagger}\\ \psi_{i-}&\psi_{i+}^{\dagger}\end{array}\right)\,,$ (27)
where $z_{i\sigma}$ are the bosonic spinons, and $\psi_{\pm}$ are charge $-e$ spinless fermions. This description has a different emergent SU(2) gauge field, but it is higgsed down to U(1) to yield the $\mathbb{CP}^{1}$ theory for the $z_{\sigma}$. The $\psi_{p}^{\dagger}$ are the holons in a holon metal carrying charges $p=\pm 1$ under the emergent U(1) [71, 72, 73, 74, 75]; see Fig. 14. The possible low $T$ fates of the holon metal have been discussed in Ref. [139].
Figure 14: Cartoon representations of the holon metal and FL* states. The holon metal has spinless charge $+e$ fermions, while FL* has spin-1/2 charge $+e$ fermions, above the same underlying spin liquid with short-range singlet valence bonds. The FL* is shown as a snapshot of the quantum dimer model of Punk et al. [124].
Numerical evidence for these dual descriptions appears in recent fuzzy sphere results [140], and it was argued [110] from these results that bosonic spinons provided a better description of the dynamic spin structure factor of this spin liquid. Which of the fermionic spinon/bosonic chargon in (20) or the bosonic spinon/fermionic chargon in (27) descriptions is more appropriate for the electronic spectrum of the thermally fluctuating pseudogap metal has not been resolved at present; indeed, both could be acceptable. Irrespective of an eventual resolution, we have argued in Section 4 that an underlying FL-FL* QPT provides at least a satisfactory description of the quantum critical fan [56].
It is worth mentioning here an important low $T$ difference between FL* and the holon metal which could be significant for future transport experiments. As the holons are spinless, the area enclosed by a holon pocket is twice that of the hole pocket in FL* [141], and the holon pocket area is the same as that of the hole pocket in the ordered antiferromagnet. The matching of ADMR to the ordered antiferromagnet Fermi surface [114] therefore supports the holon metal.
In addressing the onset of confinement at low $T$, we can exploit the duality of Wang et al. [137] to choose the formulation best suited for the confinement being considered. In the fermionic spinon description, the Higgs boson $B$ is uncondensed in the pseudogap metal, and there is a relatively straightforward method to understand confinement via the Higgs condensation of $B$. A number of puzzles on the cuprate phase diagram can be addressed by this method:
@@
The confinement crossover from the pseudogap metal to the stripy SDW state at low temperatures remains to be understood [151]. In the fermionic spinon approach, it involves a delicate interplay between the charge order associated with the condensation of $B$, and the confinement of the fermionic spinons via the SU(2) gauge field [136]. The bosonic spinon approach [73, 74, 75, 139] may be better suited to understand this crossover, and insights from numerical studies [152] should also be helpful.